# Population demographics of white dwarf binaries with intermediate separations: Gaia constraints on post-AGB mass transfer

### v1

@@ Abstract
Astrometry from the Gaia mission has revealed a large population of white dwarf (WD) + main sequence (MS) binaries with periods of $100-1000\,$d. Such systems have separations intermediate to predictions from standard binary evolution scenarios, challenging models of binary interaction and mass transfer. Because the selection function of Gaia astrometric catalogs is complex, the underlying population demographics of WD+MS binaries remain imperfectly understood. We present a forward-model of the au-scale WD+MS binary population probed by Gaia that begins with a realistic binary population and incorporates a full model of Gaia mock observations and astrometric model fitting, as well as cuts employed in producing the Gaia astrometric catalog and selecting WD+MS binary candidates. Our model allows us to constrain the intrinsic population demographics of intermediate-separation WD+MS binaries. The inferred period distribution is close to flat, with ${\rm d}N/{\rm d}P_{{\rm orb}}\propto P_{{\rm orb}}^{-0.25}$, while the WD mass distribution is sharply peaked at $0.6\,M_{\odot}$. We model the formation of au-scale WD+MS binaries as the result of interaction when the WD progenitor is an AGB star. Explaining the mass distributions of both components requires two key assumptions: (1) the mass growth of the WD is terminated when its AGB progenitor overflows its Roche lobe, and (2) post-AGB binaries binaries remain wide only if the accretor-to-donor mass ratio exceeds a critical threshold, $q_{\rm crit}\approx 0.4$. Systems with more unequal mass ratios likely shrink to shorter periods and become classical post-common envelope binaries or merge. The model implies that $\sim 1\%$ of solar-type stars have WD companions with periods of $100-1000\,$d, consistent with complementary constraints from self-lensing binaries.
@@ 1 Introduction
The abundance of these AU-scale WD+MS binaries was unexpected, because standard binary evolution models predicted few such systems to exist. While the binaries’ separations are too close to have avoided interaction when the WD progenitor was on the asymptotic giant branch (AGB) – implying they are post-MT systems – their orbits are significantly longer than predicted by models of dynamically unstable MT (i.e. common envelope evolution, CEE; e.g. Paczynski 1976; Ivanova et al. 2013). An alternative possibility is that the systems are products of stable mass transfer, but current evolutionary models predict unstable mass transfer for the systems’ inferred progenitor mass ratios (e.g. Yamaguchi et al., 2024b). Therefore, this sample highlights the imperfect current understanding of MT processes and motivates detailed studies.
@@
The rest of this paper is organized as follows: In Section 2, we summarize the steps taken by E24 to create a mock Gaia NSS catalog and describe our addition of hierarchical triples to their model. In Section 3, we provide an overview of the triage technique employed by Shahaf et al. (2024) to construct the observed sample of WD+MS binaries using orbital solutions from Gaia DR3. We also describe its application to our simulated binaries. In Section 4, we compare properties of the resulting mock sample of WD+MS binaries obtained using the original model from E24 to those of the observed sample. Based on this, in Section 5, we describe an improved model that better matches the data. In Section 6, we compare the model’s predictions to data and discuss completeness fractions and the underlying parameter distributions of WD+MS binaries. Finally, we conclude in Section 7.
@@ 2.2 Formation of compact objects
Binaries with primaries with initial masses above $8\,M_{\odot}$, which have terminated their evolution as black holes or neutron stars, are removed under the assumption that most of them are broken up or shrink to very tight orbits via CEE. Stars with initial masses below $8\,M_{\odot}$ are assumed to end their lives as WDs. E24 predicted WD masses using the empirical initial-final-mass relation (IFMR) of Weidemann (2000), which is calibrated to observations of single WDs. In this work, we explore a different IFMR that accounts for binary interaction (Section 5.2).
@@ 2.3 Post-interaction separations
E24 assumed that binaries with initial separations $<2\,$AU experience mass transfer which efficiently shrinks and circularizes their orbits such that they end up in 1-day orbits with zero eccentricities which are not detectable with astrometry. Meanwhile, they assumed that 10% of orbits with initial separations between $2$ and 666\,6AU end up in wider orbits with final separations half of their initial separations and moderate final eccentricities randomly selected between 0 and 0.2 (the remaining 90% end up in 1-day circular orbits). The 222\,2AU boundary was chosen to roughly distinguish between systems that interact with donors on the red giant branch (RGB) versus the AGB. This was motivated by recent studies of the Gaia WD+MS binary sample which have found that it is possible to from relatively wide post-common envelope binaries (PCEBs) if mass transfer begins during a thermal pulse of an AGB donor (e.g. Belloni et al., 2024; Yamaguchi et al., 2024a). These assumptions regarding the orbital evolution due to MT are modified in this work (Section 5.1).
E24 did not model the formation of low-mass helium WDs that are formed through stable mass transfer from donors on the first giant branch (e.g. Rappaport et al., 1995; Marsh et al., 1995; Brown et al., 2016; Li et al., 2019; El-Badry & Rix, 2022; Garbutt et al., 2024), and we also do not attempt to track such binaries in this work. Although low-mass WDs formed via stable mass transfer are common in samples of WD binaries selected via UV excess, they are almost entirely absent in the AMRF sample, which is only sensitive to systems with $M_{\rm WD}/M_{\star}\gtrsim 0.6$ (El-Badry, 2024).
@@ 2.4 Mock observations
We obtain mock observations for each source based on the Gaia scanning law using the Gaia Observation Forecast Tool (GOST) [^1] https://gaia.esac.esa.int/gost, and model the effect of binarity on the epoch astrometry following Lindegren (2022). We adopt an empirical noise model based on the residuals of DR3 epoch astrometry for well-behaved sources from Holl et al. (2023). Finally, we process the resulting 1D astrometric measurements through a pipeline similar to that described in Halbwachs et al. (2023) to produce a mock NSS catalog. In Figure 1, we show several simulated WD+MS binaries, along with their mock observations. The leftmost panel shows a typical WD+MS binary in the AMRF sample with $P_{\rm orb}=600\,$d at a distance of 300300300\,300pc. This enters the mock NSS catalog with a full astrometric orbital solution. In the middle panel is a case of an orbit that receives an orbital solution but is excluded from the final catalog as it does not satisfy the quality cut imposed on the relative parallax uncertainty to remove potentially spurious solutions ($\varpi/\sigma_{\varpi}<20000/P_{\rm orb}$; Halbwachs et al. 2023). Finally, the orbit on the right panel, placed at a larger distance, only receives a 7-parameter acceleration solution and therefore also does not enter the NSS catalog.
The results in E24 can be reproduced using their publicly available code [^2] https://github.com/kareemelbadry/gaiamock.
@@ 2.5 Triples
For the astrometric processing, we assume that if the angular size of the outer orbit (= semi-major axis/distance) is more than 1” on the sky, the outer tertiary is resolved and it will receive its own 5-parameter solution (Nagarajan & El-Badry, 2024). Therefore, the inner binary of such a system is processed independently. If the size of the outer orbit is less than 1” and its orbital period is less than 1000 d, the system is modeled as a binary made up of the inner binary and outer tertiary. The total mass and luminosity of the inner binary is taken as the mass and luminosity of one component. Such a system may receive an astrometric binary solution which can contaminate the AMRF sample. We discard triples with outer orbital periods greater than 1000 d with luminous tertiaries as these periods are too long for the outer orbit to be constrained in Gaia DR3, while the presence of the tertiary likely disturbs the astrometric measurements of the inner orbit (e.g. Lindegren, 2022). We also neglect systems where both components of the inner binaries have evolved into WDs, since our modeling is unlikely to reliably track the mass transfer that preceeds the second WD’s formation.
As we discuss in Section 6, the color excess cut ultimately removes the vast majority of triples, meaning that the final WD+MS binary sample is relatively insensitive to our assumptions about the triple population.
For most of the subsequent analysis, we only consider triples and WD+MS binaries, excluding all other binaries which host two luminous companions as these are almost entirely removed by the cut on the AMRF and thus do not enter the final simulated WD+MS binary sample.
In Figure 3, we show the distributions of astrometric mass function, $f_{\rm m,ast}$, of the orbital solutions in the mock NSS catalog resulting from our fiducial model (described in Section 5) as well as the true NSS catalog published in Gaia DR3. To isolate MS primaries and exclude giants, we only plot sources in the true catalog which are also present in the binary_masses catalog and consider simulated binaries with primary radii $<2\,R_{\odot}$. $f_{\rm m,ast}$ is given by
@@
It is useful for identifying both spurious astrometric solutions (Halbwachs et al., 2023) and systems with compact object companions. Here, we also include non- WD+MS binaries from the E24 whose properties are minimally affected by the changes made in this work. Looking at the simulated binaries, most of the low-$f_{\rm m,ast}$ solutions are non-WD+MS binaries while the high-$f_{\rm m,ast}$ solutions are triples. WD+MS binaries lie in between, with a median value of $0.07$. In total (not excluding giants), there are 187,000 orbital solutions in our mock catalog which is $12\%$ more than the 168,000 in the true catalog. For comparison, the original E24 catalog contained $18\%$ fewer sources than the true catalog. The difference is primarily attributed to the increased number of WD+MS binaries in our model. As we will show (Section 6), the typical WD mass in the sample is lower than assumed by E24. As a result, a smaller fraction of WD+MS binaries enter the AMRF sample than assumed in their work, and so a larger underlying WD+MS binary population is required to explain the same AMRF sample. We see that the distributions are in good agreement to one another, though it is more peaked for the mock catalog. WD + MS binaries make up $28\%$ of sources in our mock catalog, and $\sim 7\%$ of these make it into the final sample.
Figure 3: Distributions of the astrometric function, $f_{\rm m,ast}$, for the orbital solutions in the simulated fiducial (green) and true (blue) NSS catalogs. Contributions from the different components in the simulated catalog are shown with unfilled histograms. We see that the total number of orbital solutions as well as the $f_{\rm m,ast}$ distribution of the mock and true catalog are in rough agreement with each other.
@@ 3 AMRF sample of WD+MS binaries
$\mathcal{A}=\frac{\alpha}{\varpi}\left(\frac{M_{1}}{M_{\odot}}\right)^{-1/3}\left(\frac{P_{\rm orb}}{\mathrm{d}}\right)^{-2/3}$ (2)
where $\alpha$ is the angular photocentric semi-major axis, $\varpi$ is the parallax, $M_{1}$ is the primary mass, and $P_{\rm orb}$ is the orbital period. This can be rewritten in terms of the mass ratio, $q$, and G-band flux ratio, $\mathcal{S}$, of the system:
@@
For our simulated systems, we use the true primary masses from our model. This implicitly assumes that the masses reported in the binary_masses catalog are accurate and do not suffer systematic biases. We assume a fixed uncertainty of $0.06\,M_{\odot}$, which is the average value of the reported errors for the sources in the observed sample. We remove simulated binaries with primary radii $>2\,R_{\odot}$, as evolved stars are not included in the binary_masses catalog and are thus also excluded by Shahaf et al. (2024).
Using theoretical mass-luminosity relations for a range of primary masses, Shahaf et al. (2024) calculate boundaries in the $\mathcal{A}-M_{1}$ space to define several classes corresponding to the most likely types of secondaries: (1) a single MS star, (2) an inner binary of two MS stars (i.e. hierarchical triple), and (3) a dark companion. As a starting point, they isolated systems in the latter two classes (i.e. inconsistent with hosting a single MS companion), which they referred to as the non-class I sample.
For further details regarding the application of this technique to the orbits from the real NSS catalog, we refer readers to Shahaf et al. (2023, 2024).
@@ 3.1 Color excess cut
where $(B-I)_{\rm observed}$ and $(B-I)_{\rm expected}$ are the de-reddened observed and theoretically expected $B-I$ colors.
@@
Shahaf et al. (2024) took observed B, V, and I band magnitudes generated using the BP/RP spectra from the Gaia synthetic_photometry_gspc catalog and used 3D dust extinction maps (Green et al., 2019; Lallement et al., 2019) to correct for reddening. To calculate $(B-I)_{\rm expected}$, they interpolated PARSEC isochrones with a fixed age of 2 Gyrs on metallicity and the absolute V band magnitude using the stam333github.com/naamach/stam package (Hallakoun & Maoz, 2021). They obtained metallicity estimates from Zhang et al. (2023) who created an empirical forward model of the Gaia XP spectra to obtain stellar atmospheric parameters for 220 million sources. For the $\sim 20\%$ of sources flagged to have unreliable stellar parameters by Zhang et al. (2023), they set [Fe/H] = 0.0 with an uncertainty of 0.25 dex. For an in-depth description of this process, consult Section 2.2 of Shahaf et al. (2024).
@@
For each source, Shahaf et al. (2024) calculated $10^{4}$ realizations of the color excess, each time drawing observables from uncorrelated Gaussian distributions with standard deviations set to their errors. Sources where the fraction of realizations with $\Delta(B-I)>0$ exceeded 64% (corresponding to the value above which the infrared excess is greater than than 1%) were removed [^4] An error in the code used by Shahaf et al. (2024) led this threshold to be 56% in the first-published catalog. This has since been corrected and we compare our models to this updated catalog (Shahaf et al., 2025)., leaving behind $\sim 3000$ “high probability” WD+MS binaries in the final “no-color excess” (NCE) sample.
@@
For the simulated population, we predict the B, V, and I magnitudes of each system by interpolating on PARSEC isochrones and assuming the sources’ true age and metallicity. Our grid of isochrones have log(age) ranging from 6.5 to 10.5 in steps of 0.1 dex and [M/H] from -2.0 to 1.0 in steps of 0.1 dex. We calculate the observed apparent magnitudes from the total flux of all components which are then extincted/reddened according to the true distance to the source. We use the dustmaps (Green, 2018; Green et al., 2019) and mwdust python package (Bovy et al. 2016; Drimmel et al. 2003; Marshall et al. 2006; Green et al. 2019) to predict extinctions. We calculate the uncertainties in these apparent magnitudes via interpolation of an exponential fit to the flux errors against the V band fluxes of sources in the observed non-class I sample. We also fit residuals to the best-fit curve with an exponential, which we used to add noise to the interpolated errors, drawn from a normal distribution with a standard deviation given by the residual model. Finally, we convert the apparent magnitudes to “observed” absolute magnitudes using the measured parallaxes from the mock astrometric fits.
@@
We calculate $(B-I)_{\rm expected}$ using the same code as Shahaf et al. (2024), similarly assuming a fixed age of 2 Gyr. For the majority of sources, we use the true values of [Fe/H] generated by Galaxia and assume a fixed uncertainty of 0.04 dex, the median value for the observed sample. To match the observed sample, we randomly select $20\%$ of sources for which we set [Fe/H] = 0 with an elevated uncertainty of 0.25 dex.
@@
When calculating these absolute magnitudes and their uncertainties, Shahaf et al. (2024) incorrectly used parallaxes and parallax uncertainties from the gaia_source catalog, as opposed to those from the NSS catalog, which accounts for binary motion. While the parallax distribution is minimally affected by this, the reported uncertainties are, on average, a factor of 2.5 higher in the gaia_source catalog. To remain consistent with their method, we therefore increase our mock parallax uncertainty measurements by the same factor in this calculation.

### v3

@@ Abstract
Astrometry from the Gaia mission has revealed a large population of white dwarf (WD) + main sequence (MS) binaries with periods of 1001000100-1000\,d. These systems have separations intermediate to predictions from standard binary evolution scenarios, challenging models of binary interaction and mass transfer. Because the selection function of Gaia astrometric catalogs is complex, the underlying population demographics of WD+MS binaries remain imperfectly understood. We present a forward-model of the AU-scale WD+MS binary population probed by Gaia that begins with a realistic binary population and incorporates a full model of Gaia mock observations and astrometric model fitting, as well as cuts employed in producing the Gaia astrometric catalog and selecting WD+MS binary candidates. We model the formation of AU-scale WD+MS binaries as the result of interaction when the WD progenitor is an AGB star. We test several models for the binaries’ formation, including stable mass transfer with theoretically predicted stability criteria and two different formalisms for common envelope evolution. None of these models succeed in reproducing the observed component mass distributions or the absolute number of WD+MS binaries. The data are best reproduced by a model in which post-AGB binaries remain wide only if the accretor-to-donor mass ratio exceeds $\sim 0.4$. Our model allows us to constrain the intrinsic population demographics of intermediate-separation WD+MS binaries. The inferred period distribution is close to flat, with ${\rm d}N/{\rm d}P_{{\rm orb}}\propto P_{{\rm orb}}^{0.12}$, while the WD mass distribution is sharply peaked at $0.6\,M_{\odot}$. The model implies that $\sim 0.4\%$ of solar-type stars have WD companions with periods of 1001000100-1000\,d.
@@ 1 Introduction
The abundance of these AU-scale WD+MS binaries was unexpected, because standard binary evolution models predicted few such systems to exist. While the binaries’ separations are too close to have avoided interaction when the WD progenitor was on the asymptotic giant branch (AGB) – implying they are post-MT systems – their orbits are significantly longer than predicted by models of dynamically unstable MT (i.e. common envelope evolution, CEE; e.g. Paczynski 1976; Ivanova et al. 2013). An alternative possibility is that the systems are products of stable MT. While classical models predicted MT from convective donors more massive than their companions to become unstable (e.g. Paczyński, 1965; Paczyński et al., 1969; Hjellming & Webbink, 1987), recent work has shown that MT may be stable for a wider range of progenitor mass ratios (e.g. Ge et al., 2020; Temmink et al., 2023). Although several models predict MT from AGB donors to be stable for a wide range of mass ratios, they generically predict extremely high mass transfer rates ($\gtrsim 10^{-2}\,M_{\odot}\rm yr^{-1}$; e.g. Temmink et al., 2023; Yamaguchi et al., 2024a). It is not clear that binaries can survive such high mass transfer rates, since they may lead to the inflation of the accretor and ultimately, CEE (e.g. Lau et al., 2024).
Besides the objects discovered by Gaia, there are several other related classes of post-interaction binaries with comparable orbital periods, such as WD symbiotic binaries (e.g. Lü et al., 2006), post-AGB binaries (e.g. Van Winckel, 2025), and barium stars (e.g. Escorza et al., 2019), whose orbits are difficult to explain with existing binary interaction models. Insights into the mass transfer process from the Gaia WD+MS sample can also shed light on the formation history of these objects.
@@
The rest of this paper is organized as follows: In Section 2, we summarize the steps taken by E24 to create a mock Gaia NSS catalog and describe our addition of hierarchical triples to their model. In Section 3, we provide an overview of the triage technique employed by Shahaf et al. (2024) to construct the observed sample of WD+MS binaries using orbital solutions from Gaia DR3. We also describe its application to our simulated binaries. In Section 4, we compare properties of the resulting mock sample of WD+MS binaries obtained using the original model from E24 to those of the observed sample. Based on this, in Section 5, we describe several modifications to the E24 model and different models of the orbital evolution under MT to attempt to better match the data. In Section 6, we compare the models’ predictions to data and discuss completeness fractions and the underlying parameter distributions of WD+MS binaries. Finally, we conclude in Section 7.
@@ 2.2 Formation of compact objects
Binaries with primaries with initial masses above $8\,M_{\odot}$, which have terminated their evolution as black holes or neutron stars, are removed under the assumption that most of them are broken up or shrink to very tight orbits via CEE. Stars with initial masses below $8\,M_{\odot}$ are assumed to end their lives as WDs. E24 predicted WD masses using the empirical initial-final-mass relation (IFMR) of Weidemann (2000), which is calibrated to observations of single WDs. By default, we continue to use this relation in this work, but have verified that using a more updated IFMR from Cunningham et al. (2024) has a minimal effect on our final results.
@@ 2.3 Post-interaction separations
E24 assumed that binaries with initial separations <2<2\,AU experience mass transfer which efficiently shrinks and circularizes their orbits such that they end up in 1-day orbits with zero eccentricities which are not detectable with astrometry. Meanwhile, they assumed that 10% of orbits with initial separations between $2$ and 66\,AU end up in wider orbits with final separations half of their initial separations and moderate final eccentricities randomly selected between 0 and 0.2 (the remaining 90% end up in 1-day circular orbits). The 22\,AU boundary was chosen to roughly distinguish between systems that interact with donors on the red giant branch (RGB) versus the AGB. This was motivated by recent studies of the Gaia WD+MS binary sample which have found that it is possible to from relatively wide post-common envelope binaries (PCEBs) if mass transfer begins during a thermal pulse of an AGB donor (e.g. Belloni et al., 2024; Yamaguchi et al., 2024a). On the other hand, there are also recent studies showing that MT from AGB donors may remain stable for relatively unequal mass ratios, which could lead to wider post-MT orbits than expected for classical PCEBs (e.g. Ge et al., 2020; Temmink et al., 2023). These assumptions regarding the orbital evolution due to MT are modified in this work (Section 5.4).
E24 did not model the formation of low-mass helium WDs that are formed through stable mass transfer from donors on the first giant branch (e.g. Rappaport et al., 1995; Marsh et al., 1995; Brown et al., 2016; Li et al., 2019; El-Badry & Rix, 2022; Garbutt et al., 2024). Although low-mass WDs formed via stable mass transfer are common in samples of WD binaries selected via UV excess, they are almost entirely absent in the AMRF sample, which is only sensitive to systems with $M_{\rm WD}/M_{\star}\gtrsim 0.6$ (El-Badry, 2024).
@@ 2.4 Mock observations
We produce mock observations of each source using the gaiamock [^1] https://github.com/kareemelbadry/gaiamock package, which is described in detail by E24. In brief, the code produces simulated epoch astrometry following the Gaia scanning law, models the effect of binarity on the epoch astrometry following Lindegren (2022), and implements an empirical noise model based on the residuals of DR3 epoch astrometry for well-behaved sources from Holl et al. (2023). The simulated epoch astrometry is then fit with the same cascade of astrometric models used in Gaia DR3 (Halbwachs et al., 2023), ultimately producing a mock NSS catalog that can be compared to the observed NSS catalog.
In Figure 1, we show several simulated WD+MS binaries, along with their mock observations. The leftmost panel shows a typical WD+MS binary in the AMRF sample with orbital period, $P_{\rm orb}$, of 600600\,d at a distance of 300300\,pc. This system enters the mock NSS catalog with a full astrometric orbital solution. In the middle panel is a case of an orbit that receives an orbital solution but is excluded from the final catalog because it does not pass a quality cut imposed on the parallax signal-to-noise ratio to remove spurious solutions ($\varpi/\sigma_{\varpi}>20000/P_{\rm orb}$; Halbwachs et al. 2023). Finally, the orbit shown in the right panel, placed at a larger distance, only receives a 7-parameter acceleration solution and therefore also does not enter the NSS catalog.
@@ 2.5 Triples
For the astrometric processing, we assume that if the angular size of the outer orbit (= semi-major axis/distance) is more than 1” on the sky, the outer tertiary is resolved and will receive its own 5-parameter solution (Nagarajan & El-Badry, 2024). Therefore, the inner binary of such a system is processed independently. If the size of the outer orbit is smaller than 1” and its orbital period is less than 1000 d, the system is modeled as a binary made up of the inner binary and outer tertiary. The total mass and luminosity of the inner binary is taken as the mass and luminosity of one component. Such a system may receive an astrometric binary solution that can contaminate the AMRF sample. We discard triples with outer orbital periods longer than 1000 d and luminous tertiaries: these periods are too long for the outer orbit to be constrained in Gaia DR3, and the presence of the tertiary will disturb the astrometric measurements of the inner orbit (e.g. Lindegren, 2022). We also neglect systems where both components of the inner binaries have evolved into WDs, since our modeling is unlikely to reliably track the mass transfer that preceeds the second WD’s formation.
As we discuss in Section 6, the color excess cut ultimately removes the vast majority of triples, meaning that the final WD+MS binary sample is relatively insensitive to our assumptions about the triple population. However, modeling of triples is important for interpretation of the AMRF sample, because some genuine WD+MS binaries are excluded by cuts designed to filter out triples.
For most of the subsequent analysis, we only consider triples and WD+MS binaries, excluding all other binaries which host two luminous companions. The latter are almost entirely removed by cuts on the AMRF and thus do not enter the final simulated WD+MS binary sample.
In Figure 3, we show the distributions of astrometric mass function, $f_{\rm m,ast}$, of the orbital solutions in the mock NSS catalog resulting from our empirical model (described in Section 5.4.4) as well as the true NSS catalog published in Gaia DR3. To isolate MS primaries and exclude giants, we only plot sources in the observed catalog which are also present in the binary_masses catalog and only consider simulated binaries with primary radii $2\,R_{\odot}$. $f_{\rm m,ast}$ is given by
@@
This quantity is useful for identifying both spurious astrometric solutions (Halbwachs et al., 2023) and systems with compact object companions.
In the simulated sample, most of the low-$f_{\rm m,ast}$ solutions are non-WD+MS binaries, while the highest-$f_{\rm m,ast}$ solutions are triples. WD+MS binaries lie in between these populations, with a median $f_{\rm m,ast}$ of $0.09$. In total (not excluding giants), there are 156,441 orbital solutions in the mock catalog, which is $7\%$ fewer than the 168,000 in the observed catalog. For comparison, the original E24 catalog contained $18\%$ fewer sources than the true catalog. The difference is primarily due to the larger number of WD+MS binaries in our model. As we will show (Section 6), the typical WD mass in the sample is lower than assumed by E24. As a result, a smaller fraction of WD+MS binaries enter the AMRF sample than assumed in their work, and so a larger underlying WD+MS binary population is required to explain the same AMRF sample. The observed and simulated $f_{\rm m,ast}$ distributions are in reasonably good agreement. WD+MS binaries make up $14\%$ of sources in the mock catalog, and $15\%$ of these make it into the final WD+MS sample. This implies that the observed NSS catalog likely contains many WD+MS binaries that have yet to be identified.
Figure 3: Distributions of the astrometric function, $f_{\rm m,ast}$, for the orbital solutions in the simulated empirical (green) and true (blue) NSS catalogs. Contributions from the different components in the simulated catalog are shown with unfilled histograms. We see that the total number of orbital solutions as well as the $f_{\rm m,ast}$ distribution of the mock and true catalog are in rough agreement with each other.
@@ 3 AMRF sample of WD+MS binaries
$\mathcal{A}=\frac{\alpha}{\varpi}\left(\frac{M_{1}}{M_{\odot}}\right)^{-1/3}\left(\frac{P_{\rm orb}}{\mathrm{yr}}\right)^{-2/3}$ (2)
where $\alpha$ is the angular photocentric semi-major axis, $\varpi$ is the parallax, $M_{1}$ is the primary mass, and $P_{\rm orb}$ is the orbital period. The AMFRF can be rewritten in terms of the mass ratio, $q$, and G-band flux ratio, $\mathcal{S}$, of the system:
@@
For our simulated systems, we use the true primary masses from our model, implicitly assuming that the masses reported in the binary_masses catalog are accurate and do not suffer systematic biases. We assume a fixed uncertainty of $0.06\,M_{\odot}$ in $M_{1}$, which is the average uncertainty in the observed sample. We remove simulated binaries with primary radii $2\,R_{\odot}$, as evolved stars are not included in the binary_masses catalog and are thus also excluded by Shahaf et al. (2024).
Using theoretical mass-luminosity relations for a range of primary masses, Shahaf et al. (2024) calculate boundaries in the $\mathcal{A}-M_{1}$ space to define several classes corresponding to the most likely types of secondaries: (1) a single MS star, (2) an inner binary of two MS stars (i.e. hierarchical triple), and (3) a dark companion. As a starting point, they isolated systems in the latter two classes (i.e. inconsistent with hosting a single MS companion), which they call the “non-class I” sample. For details regarding the application of this technique to the orbits from the real NSS catalog, we refer readers to Shahaf et al. (2023, 2024).
@@ 3.1 Color excess cut
where $(B-I)_{\rm observed}$ and $(B-I)_{\rm expected}$ are the de-reddened observed and theoretically expected BIB-I colors.
@@
Shahaf et al. (2024) used observed B, V, and I band magnitudes generated using the BP/RP spectra from the Gaia synthetic_photometry_gspc catalog and used 3D dust extinction maps (Green et al., 2019; Lallement et al., 2019) to correct for reddening. To calculate $(B-I)_{\rm expected}$, they interpolated PARSEC isochrones with a fixed age of 2 Gyrs on metallicity and the absolute V band magnitude using the stam222github.com/naamach/stam package (Hallakoun & Maoz, 2021). They assumed a fixed age of 2 Gyr because the ages of the observed systems are unknown. In Section 3.1.1, we discuss the effect of this assumption on the simulated sample. They obtained metallicity estimates from Zhang et al. (2023) who created an empirical forward model of the Gaia XP spectra to obtain stellar atmospheric parameters for 220 million sources. For the $\sim 20\%$ of sources flagged by Zhang et al. (2023) to have unreliable stellar parameters, Shahaf et al. (2024) set [Fe/H] = 0.0, with an uncertainty of 0.25 dex.
@@
For each source, Shahaf et al. (2024) calculated $10^{4}$ samples of the color excess, each time drawing observables from uncorrelated Gaussian distributions with standard deviations set to their errors. Sources where the fraction of realizations with $\Delta(B-I)>0$ exceeded 64% (corresponding to the value above which the infrared excess is greater than than 1%) were removed [^3] An error in the code used by Shahaf et al. (2024) led this threshold to be 56% in the first-published catalog. This has since been corrected, and we compare our models to this updated catalog (Shahaf et al., 2025)., leaving behind $\sim 3000$ high probability WD+MS binaries in the final “no-color excess” (NCE) sample.
@@
For the simulated population, we predict the B, V, and I magnitudes of each system by interpolating on PARSEC isochrones and assuming the sources’ true age and metallicity. Our grid of isochrones has log(age) ranging from 6.5 to 10.5 in steps of 0.1 dex and [M/H] from -2.0 to 1.0 in steps of 0.1 dex. We calculate the observed apparent magnitudes from the total flux of all components, which are then extincted/reddened according to the true distance to the source. We use the dustmaps (Green, 2018; Green et al., 2019) and mwdust python package (Bovy et al. 2016; Drimmel et al. 2003; Marshall et al. 2006; Green et al. 2019) to predict extinctions. We calculate the uncertainties in these apparent magnitudes via interpolation of an exponential fit to the flux errors as a function of the V band fluxes of sources in the observed non-class I sample. We also fit residuals to the best-fit curve with an exponential, which we used to add noise to the interpolated errors, drawn from a normal distribution with a standard deviation given by the residual model. Finally, we convert the apparent magnitudes to “observed” absolute magnitudes using the measured parallaxes from the mock astrometric fits.
@@
We calculate $(B-I)_{\rm expected}$ using the same code as Shahaf et al. (2024), similarly assuming a fixed age of 2 Gyr. For the majority of sources, we use the true values of [Fe/H] generated by Galaxia and assume a fixed uncertainty of 0.04 dex, the median value for the observed sample. To match the observed sample, we randomly select $20\%$ of sources to have unreliable metallcities; for these, we set [Fe/H] = 0 and an elevated uncertainty of 0.25 dex.
@@
When calculating these absolute magnitudes and their uncertainties, Shahaf et al. (2024) incorrectly used parallaxes and parallax uncertainties from the gaia_source catalog, as opposed to those from the NSS catalog, which accounts for binary motion. While the parallax distribution is minimally affected by this, the reported uncertainties are, on average, a factor of 2.5 larger in the gaia_source catalog. To remain consistent with their method, we therefore increase our mock parallax uncertainty measurements by the same factor in this calculation.